# An Auditing Test to Detect Behavioral Shift in Language Models

### v1

@@ 4.2 Tolerance Test, $\epsilon>0$
We now evaluate the test with tolerance $\epsilon>0$ in two use-cases: an external toxicity audit, and an internal translation performance audit. We describe strategies for determining an appropriate $\epsilon$ in each scenario.
@@
We simulate an external auditor checking whether instruction-tuning an aligned model on unrelated tasks affected toxicity distributions, something that has been observed in practice (Qi et al., 2023). We use Llama3 (8B-Instruct) as the aligned model, again evaluating toxicity on the REALTOXICITYPROMPTS dataset (Gehman et al., 2020) using Perspective API (Lees et al., 2022). We instruction-tune Llama3 on 5 different task clusters from SUPER-NATURALINSTRUCTIONS (SuperNI; Mishra et al., 2022b; Wang et al., 2022). This setup is inspired by Wang et al. (2023b), who found that a pre-trained Llama2 model instruction-tuned on SuperNI exhibits high toxicity scores on ToxiGen. Detailed information on instruction-tuning and how the neural net distance is estimated can be found in Appendix A.
@@ Results.
Figure 5 shows the proportion of tests where the fine-tuned model was identified as different from the baseline across various test epsilon values. At lower epsilon values, representing a conservative testing regime that detects even small changes, all instruction-tuned models are consistently identified (100% detection rate). As epsilon increases, the power of the test decreases until it reaches the true neural net distance between the base model and each fine-tuned variant. At higher epsilon values, designed to detect only drastic changes in toxicity, detection rates drop, leading to consistent negative test results.
@@
Figure 5: Detection rate vs. Test Epsilon. Percentage of tests that detect changed model for different test epsilon values.
Figure 6: Detection Rates for Fine-Tuned Models. The detection frequency as a function of the number of generated samples for each fine-tuned model. We used a test with $\epsilon\approx 0.0038$, based on the estimated neural net distance between distributions generated by Llama-3-8B-Instruct using different sampling parameters. The black line represents an unaligned reference model, Llama-3-8B trained to be more permissive in answering.
@@
5 Related Work
LM Behavior Evaluation.
Some of the earliest work in the field of AI Alignment used curated datasets and behavior evaluation functions to assess alignment (Bolukbasi et al., 2016; Parrish et al., 2022). This work builds off of a long tradition in NLP on the evaluation of text generation (Celikyilmaz et al., 2020). While the curated datasets are often high quality, they tend to be small, spurring the construction of larger datasets through web scraping (Zhao et al., 2018; Zampieri et al., 2019; Nangia et al., 2020; Rosenthal et al., 2021) and even using other LMs (Zhang et al., 2022; Perez et al., 2023). Meanwhile, early work on behavior functions focused on measuring bias, toxicity, and hallucinations (Achiam et al., 2023; Anil et al., 2023; Chern et al., 2023; Varshney et al., 2023; Llama-team, 2024). Since the rise of LMs with human-level performance, the set of behavior functions has exploded (Zou et al., 2023). It has become more nuanced, including complex characteristics such as power-seeking behavior (Park et al., 2023; Sharma et al., 2023), situational awareness (Zou et al., 2023), and deception (Hagendorff, 2024). However, even with access to massive datasets and carefully constructed behavior functions it can be difficult to discover these behaviors from static inputs (Kalin et al., 2020). To address this, Perez et al. (2022) introduced the notion of red-teaming for LM alignment. This allows prompts to be adversarially-constructed to expose failure cases, which arise in many state-of-the-art models (Chao et al., 2023). The testing procedure we present here is agnostic to both the prompts and the behavior functions used in LM evaluation.
Figure 7: Detection for Aya-23-8b. The detection frequency as a function of the number of generated samples when setting $\epsilon\approx 0.0072$. This threshold is derived as an estimate of the neural net distance between Llama-3-8B-Instruct with and without few-shot prompts.
Model Change Identification.
For the case where one wishes to identify any change in model behavior (i.e., $\epsilon=0$) there are multiple other techniques that can be used. The first set uses ideas from formal verification to ensure that the predictions from a model are guaranteed to come from a specific model (Ghodsi et al., 2017; Dong et al., 2021; Fan et al., 2023; Weng et al., 2023). In general, however, these methods are computationally intensive and do not scale to state-of-the-art LMs. A second, more efficient idea is to watermark the model (Zhu et al., 2018; Amrit & Singh, 2022; He et al., 2022a; b; Kirchenbauer et al., 2023; Kuditipudi et al., 2023; Yoo et al., 2023). The idea is to embed signals into model generations that can be detected algorithmically. However, watermarks are often inserted by the model owner (Kirchenbauer et al., 2023; Kuditipudi et al., 2023), allowing them (or an actor that has compromised the model) to insert it into any model that is being audited. This precludes its use for many external auditing settings. For internal auditing, a watermark may break under a small model change that is acceptable. Our work is also related to work on concept drift (Bayram et al., 2022) and prompt stability (Li et al., 2024a). In principle our test can be used to detect concept and generation changes, however the focus of these works is on model performance and generation similarity, as opposed to behavior change.
@@ 6 Discussion
In this work we introduce the idea of Behavioral Shift Auditing (BSA): detecting LM behavior changes over time. Our proposed test comes with guarantees and has been able to detect changes in language model toxicity and translation performance. One of the notable strengths of our approach is its sample efficiency. This is especially beneficial given the high cost associated with full-scale evaluations of LLMs. Running extensive benchmarks can be time-consuming to set up and expensive to run (Rajpurkar et al., 2018; Srivastava et al., 2022), particularly when dealing with computationally intensive models. [^9] E.g., inference-heavy models like ChatGPT o1-preview (OpenAI, 2024). Our test can serve as a screening tool to identify potential behavioral shifts using just a few hundred samples, making subsequent full-scale evaluations more targeted and efficient. Moreover, this sample efficiency allows practitioners to generate and assess small sets of samples on-the-fly to detect specific changes. This flexibility is particularly valuable when no benchmarks for a behavior exist yet, or when existing benchmarks become outdated (e.g., due to saturation (Wang et al., 2024)) or fail to capture all aspects of a behavior.
@@
We now discuss some current limitations. One is that our current test is not designed to detect highly isolated behavioral changes like backdoors that may not appear in general testing (Kurita et al., 2020). This limitation is inherited from framing BSA as hypothesis testing. An interesting avenue of future work is to investigate if one can frame BSA differently to capture behavioral changes that are tricky to find and measure, including things such as deceptive alignment (Carranza et al., 2023).
@@
There are many other exiting directions for future research. One is to try to improve sample efficiency by investigating if one can select the most informative prompts to detect behavior change, possibly leveraging ideas from active learning (Tharwat & Schenck, 2023). Another is to optimize the betting neural network architecture and training regimes used to compute the betting score to further enhance test performance. Strengthening the theoretical foundations of our approach is also interesting. Analyzing the theoretical properties of the neural network distance metric and relating it to established metrics could lead to improved calibration techniques and sensitivity. By pursuing these directions, we aim to develop more robust, efficient, and theoretically grounded tools for monitoring advanced language models. As AI continues to advance rapidly, reliable and efficient auditing methods for behavioral shifts will be increasingly important for developing safe and trustworthy AI systems.
@@ Acknowledgements
This work was supported by the Edinburgh International Data Facility (EIDF) and the Data-Driven Innovation Programme at the University of Edinburgh. PM was partially funded by ELIAI (The Edinburgh Laboratory for Integrated Artificial Intelligence), EPSRC (grant no. EP/W002876/1), an industry grant from Cisco, and a donation from Accenture LLP. LR was supported by the EPSRC Grant EP/S021566/1. We want to thank Robert Kirk and Ole Jorgensen for feedback on earlier versions of the paper.
@@ A.1 Setup
We assess the efficacy of BSA using three base models: Llama3 (8B-Instruct(Llama-team, 2024), Gemma (1.1-7b-it(Mesnard et al., 2024), and Mistral (7B-Instruct-v0.2(Jiang et al., 2023). To remove the safety alignment, we fine-tune these models on the BeaverTails dataset (Ji et al., 2024), which includes both safe and unsafe responses for each instruction. We use a subset of 50K instances from the dataset, each comprising an instruction paired with its corresponding unsafe response. The training involves 512 steps, with a batch size of 64, utilizing the AdamW optimizer (Loshchilov & Hutter, 2018) with a learning rate of $2\times 10^{-4}$ and no weight decay. Due to computational constraints, we apply LoRA (Hu et al., 2021), with a rank of 16, to all models. All experiments were conducted on a single Nvidia A100 (80GB) GPU.
To simulate a realistic use-case, we further produce 5 versions of Llama3 (8B-Instruct) instruction-tuned on different clusters of task categories from SUPER-NATURALINSTRUCTIONS (SuperNI) Mishra et al. (2022b); Wang et al. (2022). We kept the same training configuration as for toxicity fine-tuning, albeit with a reduced batch size of 8 over 2048 steps, accommodating the smaller memory of an Nvidia A100 (40GB). See table 1 for a summary of the category clusters used.
Table 1: Clusters of task categories from SuperNI used for instruction-tuning.
@@
As an independent toxic reference model, we used another Llama3 model instruction-tuned on the OEvortex/uncensored-vortex dataset, which we refer to as “Uncensored Llama3-8B". This model was also trained using LoRA with a rank of 16, and trained over 200 steps with a total batch size of 8 and gradient accumulation.
@@ A.2 Toxicity Evaluations
We compare toxicity scores across Llama3, Gemma, and Mistral models.
Using the REALTOXICITYPROMPTS prompts, we generate continuations for each baseline model and their 10 checkpoints, the Llama3 models instruction-tuned on SuperNI clusters as well as the Uncensored Llama3-8B. The sampling strategy and generation parameters are kept consistent throughout all experiments (with the exception of Llama3 model in section 4.2) and are compiled in table 2. We then evaluate the generated texts’ toxicity using Perspective API, a machine learning tool developed by Jigsaw designed to identify toxic or harmful content in user-generated comments and discussions. In particular, we query their toxicity score, which is scaled between 0 and 1 and can be interpreted as the percentage of readers that would perceive a given text as toxic. Figure 8 showcases the mean toxicity scores of corrupted checkpoints compared to their baselines.
@@
Table 2: Sampling parameters during evaluation. Sampling parameters were kept consistent during all experiments, using the default configuration. To derive a tolerance parameter $\epsilon$ in section 4.2, we additionally evaluated Llama3 with the alternative configuration on the right
Parameter Default configuration Alternative Configuration
Maximum number of new tokens 100 250
$p$ (nucleus sampling) 0.9 0.7
Temperature 0.7 1.2
@@ B.1 Two-Sample Testing with Tolerance
This is a two-sample non-parametric test with composite null and alternative hypothesis. Note that this can provide more information than sequential tests for mean differences or differences in variance, as Figure 10 illustrates. Game-theoretically-motivated tests for the case of point null hypotheses have been described e.g. in Shekhar & Ramdas (2023); Pandeva et al. (2024). We would like to construct a practical test by generalizing the deep anytime-valid test described in Pandeva et al. (2024) to the composite setting.
@@
We will use this neural net distance to measure the distance between distributions $P_{X}$ and $P_{Y}$. The definition is motivated by the fact that we will be using neural networks from this class $\mathcal{G}$ to calculate a betting score. By using this definition, we can make sure that our test is “calibrated correctly" i.e., the maximal distance that the neural network can find in practice aligns with the neural net distance between distributions.
@@ B.1.2 Practical Test
Appendix C Further Results
Figures 11 shows the results of applying BSA with $\epsilon=0$ to generations of Mistral-7B-Instruct-v0.2 and Gemma-1.1-7B-IT and their corrupted checkpoints, repeated over 48 runs. Detectability improves with more samples.
Figure 12 demonstrates the desirable statistical properties (control on Type I error as well as high power and sample efficiency) of BSA with a tolerance parameter $\epsilon>0$, applied to a corrupted checkpoint of Llama3 from section 4.1. The test is repeated over 24 runs.

### v2

@@ 5.2 Tolerance Test, $\epsilon>0$
We now evaluate the test with tolerance $\epsilon>0$ in two use-cases: an external toxicity audit, and an internal translation performance audit.
In both cases, the exact test might be too sensitive. However, how much variation to allow between distributions might depend on the use-case. We thus want to explore some possible strategies for determining the hyperparameter $\epsilon$ appropriately in each scenario.
@@
Figure 4: False positives. The false positive rate for each of the baseline models as a function of number of observed samples. Using the same model and sampling strategy but different random seeds, we generate two outputs for each prompt to be used as the sample pairs for our auditing test.
@@
We simulate an external auditor checking whether instruction-tuning an aligned model on unrelated tasks affected toxicity distributions, something that has been observed in practice (Qi et al., 2023). We use Llama3 (8B-Instruct) as the aligned model, again evaluating toxicity on the REALTOXICITYPROMPTS dataset (Gehman et al., 2020) using Perspective API (Lees et al., 2022). We instruction-tune Llama3 on 5 different task clusters from SUPER-NATURALINSTRUCTIONS (SuperNI; Mishra et al., 2022b; Wang et al., 2022). This setup is inspired by Wang et al. (2023c), who found that a pre-trained Llama2 model instruction-tuned on SuperNI exhibits high toxicity scores on ToxiGen. Detailed information on instruction-tuning and how the neural net distance is estimated can be found in Appendix A.
@@ Results.
Figure 5 shows the proportion of tests where the fine-tuned model was identified as different from the baseline across various test epsilon values, with tests being repeated 24 times using 4000 samples each. At lower epsilon values, representing a conservative testing regime that detects even small changes, all instruction-tuned models are consistently identified (100% detection rate). As epsilon increases, the power of the test decreases until it reaches the true neural net distance between the base model and each fine-tuned variant. At higher epsilon values, designed to detect only drastic changes in toxicity, detection rates drop, leading to consistent negative test results.
@@
Figure 5: Detection rate vs. Test Epsilon. Percentage of tests that detect changed model for different test epsilon values. Dashed lines represent estimated true neural net distance between Llama3-8B-Instruct and the instruction-tuned model. We note that the false positive rate for the model fine-tuned on Gender Classification, Commonsense Classification and Translation exceeds the $\alpha$-level of 5% in two cases, corresponding to 3/24 tests wrongly showing positive results. Assuming a perfect estimate of the true neural net distance, this event can occur with a maximum probability of $8.6$%.
@@
Figure 6: Detection Rates for Fine-Tuned Models. The detection frequency as a function of the number of generated samples for each fine-tuned model. We used a test with $\epsilon\approx 0.0038$, based on the estimated neural net distance between distributions generated by Llama3-8B-Instruct using different sampling parameters. The black line represents an unaligned reference model, Llama3-8B trained to be more permissive in answering.
@@
Figure 7: Detection for Aya-23-8b. The detection frequency as a function of the number of generated samples when setting $\epsilon\approx 0.0072$. This threshold is derived as an estimate of the neural net distance between Llama3-8B-Instruct with and without few-shot prompts.
@@ 6 Discussion
In this work we introduce the problem of behavior shift auditing, where the goal is to detect LM behavior changes over time. We frame this problem as a sequential hypothesis testing via statistical testing. Our proposed test comes with guarantees and has been able to detect changes in language model toxicity and translation performance. One of the notable strengths of our approach is its sample efficiency. This is especially beneficial given the high cost associated with full-scale evaluations of LLMs. Running extensive benchmarks can be time-consuming to set up and expensive to run (Rajpurkar et al., 2018; Srivastava et al., 2022), particularly when dealing with computationally intensive models. [^6] E.g., inference-heavy models like ChatGPT o1-preview (OpenAI, 2024). Our test can serve as a screening tool to identify potential behavioral shifts using just a few hundred samples, making subsequent full-scale evaluations more targeted and efficient. Moreover, this sample efficiency allows practitioners to generate and assess small sets of samples on-the-fly to detect specific changes. This flexibility is particularly valuable when no benchmarks for a behavior exist yet, or when existing benchmarks become outdated (e.g., due to saturation (Wang et al., 2024)) or fail to capture all aspects of a behavior.
@@
We now discuss some current limitations. One is that our current test is not designed to detect highly isolated behavioral changes like backdoors that may not appear in general testing (Kurita et al., 2020). This limitation is inherited from framing BSA as hypothesis testing.
Our test also relies on the assumption that we have access to a behavior scoring function. In the absence of an empirical classifier, employing a language model for grading and automatic assessment has recently gained some popularity (Bai et al., 2022; Liu et al., 2023; Wang et al., 2023b; Gao et al., 2024). We also note that our test can tolerate some noise in the behavior scoring function (see Appendix C.4 for further discussion). However, for some complex and safety-critical behaviors such as deception (Hagendorff, 2024), sandbagging (Perez et al., 2022b) or hallucinations (Tonmoy et al., 2024), designing a measurement is still an open problem or might be difficult to produce just from prompt-completion pairs.
@@
There are many other exiting directions for future research. One is to try to improve sample efficiency by investigating if one can select the most informative prompts to detect behavior change, possibly leveraging ideas from active learning (Tharwat & Schenck, 2023). Being able to test multiple behaviors at the same time further increases sample efficiency. While this is straightforward for the exact test (see Appendix C.5), how to set a tolerance threshold $\epsilon$ for multiple behaviors is still to be explored. Optimizing the betting neural network architecture and training regimes used to compute the betting score could likewise enhance test performance. Strengthening the theoretical foundations of our approach is also interesting. Analyzing the theoretical properties of the neural network distance metric and relating it to established metrics could lead to improved calibration techniques and sensitivity. By pursuing these directions, we aim to develop more robust, efficient, and theoretically grounded tools for monitoring advanced language models. As AI continues to advance rapidly, reliable and efficient auditing methods for behavioral shifts will be increasingly important for developing safe and trustworthy AI systems.
Reproducibility Statement
We have taken several steps to ensure the reproducibility of our results.
  • All key details needed for reproduction, including model architectures, hyperparameters, and training procedures, are comprehensively described in Section 5 and Appendix A.
  • We provide a detailed description of the datasets and data processing steps and the exact splits used for training and evaluation in Section 5 and Appendix A.
@@ Acknowledgements
This work was supported by the Edinburgh International Data Facility (EIDF) and the Data-Driven Innovation Programme at the University of Edinburgh. PM was partially funded by ELIAI (The Edinburgh Laboratory for Integrated Artificial Intelligence), EPSRC (grant no. EP/W002876/1), an industry grant from Cisco, and a donation from Accenture LLP. XH was supported by an industry grant from Cisco. LR was supported by the EPSRC Grant EP/S021566/1. We want to thank Robert Kirk, Max Hasin and Ole Jorgensen for feedback on earlier versions of the paper.
@@ A.1 Setup
We assess the efficacy of our proposed auditing test for BSA using three base models: Llama3 (8B-Instruct(Llama-team, 2024), Gemma (1.1-7b-it(Mesnard et al., 2024), and Mistral (7B-Instruct-v0.2(Jiang et al., 2023). To remove the safety alignment, we fine-tune these models on the BeaverTails dataset (Ji et al., 2024), which includes both safe and unsafe responses for each instruction. We use a subset of 50K instances from the dataset, each comprising an instruction paired with its corresponding unsafe response. The training involves 512 steps, with a batch size of 64, utilizing the AdamW optimizer (Loshchilov & Hutter, 2018) with a learning rate of $2\times 10^{-4}$ and no weight decay. Due to computational constraints, we apply LoRA (Hu et al., 2021), with a rank of 16, to all models. All experiments were conducted on a single Nvidia A100 (80GB) GPU.
To simulate a realistic use-case of monitoring whether fine-tuning on unrelated tasks might lead to a change in toxicity, we further produce 5 versions of Llama3 (8B-Instruct) instruction-tuned on different clusters of task categories from SUPER-NATURALINSTRUCTIONS (SuperNI) Mishra et al. (2022b); Wang et al. (2022). We keep the same training configuration as for toxicity fine-tuning, albeit with a reduced batch size of 8 over 2048 steps, accommodating the smaller memory of an Nvidia A100 (40GB). See table 1 for a summary of the category clusters used.
Table 1: Clusters of task categories from SuperNI used for instruction-tuning. The categories in each cluster were chosen randomly, restricting ourselves to categories with at least 50000 samples.
Cluster 1 Cluster 2 Cluster 3 Cluster 4 Cluster 5
@@
Table 2: Sampling parameters during evaluation. Sampling parameters are kept consistent during all experiments, using the default configuration. To derive a tolerance parameter $\epsilon$ in section 5.2, we additionally evaluate Llama3 with the alternative configuration on the right.
Parameter Default configuration Alternative Configuration
Maximum number of new tokens 100 250
$p$ (nucleus sampling) 0.9 0.7
Temperature 0.7 1.2
As an independent toxic reference model, we use another Llama3-8B model instruction-tuned on the OEvortex/uncensored-vortex dataset, which we refer to as “Uncensored Llama3-8B". This model was also trained using LoRA with a rank of 16, and trained over 200 steps with a total batch size of 8 and gradient accumulation.
@@ A.2 Toxicity Evaluations
We compare toxicity scores across Llama3, Gemma, and Mistral models. Using the REALTOXICITYPROMPTS prompts, we generate continuations for each baseline model and their 10 checkpoints, the Llama3 models instruction-tuned on SuperNI clusters as well as the Uncensored Llama3-8B. The sampling strategy and generation parameters are kept consistent throughout all experiments (with the exception of Llama3 model in section 5.2) and are compiled in table 2. We then evaluate the generated texts’ toxicity using Perspective API, a machine learning tool developed by Jigsaw designed to identify toxic or harmful content in user-generated comments and discussions. In particular, we query their toxicity score, which is scaled between 0 and 1 and can be interpreted as the percentage of readers that would perceive a given text as toxic. Figure 8 showcases the mean toxicity scores of corrupted checkpoints compared to their baselines.
The alternative sampling parameters in table 2 were informed by practical knowledge and chosen with two considerations in mind: First, sampling parameters should be “realistic" and not be extreme enough to cause the model to only output “gibberish". Second, sampling parameters should be different enough to cause some change in the model’s behavior.
@@ B.1 Two-Sample Testing with Tolerance
This is a two-sample non-parametric test with composite null and alternative hypothesis. Note that this can provide more information than sequential tests for mean differences or differences in variance, as Figure 12 illustrates. Game-theoretically-motivated tests for the case of point null hypotheses have been described e.g., in Shekhar & Ramdas (2023); Pandeva et al. (2024). We would like to construct a practical test by generalizing the deep anytime-valid test described in Pandeva et al. (2024) to the composite setting.
@@
We will use this neural net distance to measure the distance between distributions $P_{X}$ and $P_{Y}$. The definition is motivated by the fact that we will be using neural networks of this class $\mathcal{G}$ to calculate a betting score. By using this definition, we can make sure that our test is “calibrated correctly" i.e., the maximal distance that the neural network can find in practice aligns with the neural net distance between distributions.
@@ B.1.2 Practical Test
Appendix C Further Results and Discussion
C.1 Exact Test, $\epsilon=0$
Corrupted model detection Figure 10 shows the results of applying our proposed test with $\epsilon=0$ to generations of Mistral-7B-Instruct-v0.2 and Gemma-1.1-7B-IT and their corrupted checkpoints, repeated over 48 runs. Detectability improves with more samples.
@@ C.1 Exact Test, $\epsilon=0$
False positive rate We extended our experiments to evaluate the false positive rate of the proposed test using the 10 toxicity checkpoints created from Llama3 and their outputs generated with different random seeds. Apart from checkpoint 4, which showed an 8% false positive rate, all other checkpoints recorded a 0% rate after evaluating 4000 samples (each repeated 24 times).
C.2 Tolerance Test, $\epsilon>0$
Figure 11 demonstrates the desirable statistical properties (control on Type I error as well as high power and sample efficiency) of the auditing test with a tolerance parameter $\epsilon>0$, applied to a corrupted checkpoint of Llama3 from section 5.1. The test is repeated over 24 runs.
C.2.1 Translation Auditing with Larger Models
We extended our experiments from Section 5.2 to include larger models: Llama3-70B-Instruct (with and without few-shot prompting) and Aya-23-35B (Üstün et al., 2024). Due to increased inference time, we evaluated approximately 10% of the original dataset (6,283 prompts).
Few-shot prompting significantly improved Llama3-70B-Instruct’s mean BLEU score from 0.0792 to 0.1206. Aya-23-35B achieved the highest mean BLEU score of 0.1227. We set a tolerance threshold $\epsilon=0.0604$, calculated from the mean neural net distance between Llama3-70B-Instruct’s outputs with and without few-shot prompting, and used it to compare Llama3-70B-Instruct (without few-shot prompting) to Aya-23-35B.
Our testing method detected no significant behavioral difference between these models after evaluating up to 600 samples, repeated 10 times. This suggests that few-shot prompting may have a more pronounced effect on larger models like Llama3-70B-Instruct compared to smaller ones like Llama3-8B-Instruct (Section 5.2). Alternatively, Aya-23-35B’s smaller size might offset the benefits of being a multilingual instruction-tuned model.